Note: (show all your works and measure that you answer all small questions from part A to part B below. Sorry for long questions, but I give thanks to anyone who can help me)A certain reaction has a. Is the process spontaneous at 238K equilibrium consistent with the free energy b. What is the equilibrium constant for the reaction at 298K? Is the magnitude of the calculated? Explain. B. A voltaic cell is made with Mn/Mn Cd/Cd t. It has the following initial concentrations .060M and Cd2 090M cell 78 V b. is value of Ecol when the concentrations ofions are equal? of that cell can do? c. What is the value of Ecell and AG when equilibrium is reached? How does this relate to the amount work the
Answer #1

Solved the first problem, post multiple question to get the remaining answer

A)

a) The reaction is spontaneous, if the Gibbs free energy is less than zero

\Delta G = \Delta H - T \Delta S

\Delta G = 15400 - (64.7)(238) = 1.4 J/mol

Since the value of Delta G is positive at 238K, hence the reaction is not spontaneous at that temperature

b)

\Delta G = \Delta H - T \Delta S

\Delta G = 15400 - (64.7)(298) = -3880.6 J/mol

\Delta G = -RTln(K)

3880.6 = (8.314)(298)ln(K)

ln(K) = 1.5662

K = e^{1.5662} = 4.7884